OpenStudy (anonymous):

Easy question I know. I think the answer is 5 but I need someone to check if its right and if it isnt explain please. I need to get an A on this. (Medal + Fan) ASAP If y varies inversely as x and the constant of variation is thirty, what is y when x equals six? a) 5 b) 24 c) 36 d) 180

OpenStudy (anonymous):

y = 5

OpenStudy (anonymous):

if y varies inversely as x: y = k/x, k is the constant of variation y = 30/6 y = 5

OpenStudy (anonymous):

@Peasant1 haha I knew it, Im was just asking I needed a check up.. Thanks! I feel so stupid lol

OpenStudy (anonymous):

@Juliette32801, 1. Because you needed help doesn't make you stupid. I need clarification all the time. :P 2. No problem. Feel free to ask anytime. I'll help if I can! :D

OpenStudy (anonymous):

@Peasant1 Thanks for understanding :))

OpenStudy (anonymous):

No problem! :D